# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v3

@@ Abstract
I Introduction
@@
Proper performance analysis of devices is key to provide information to potentially interested users, allowing fair comparisons among different devices. Additionally, manufacturers use these insights to keep improving their systems. The measurement of performance for classical computer systems has traditionally been controversial. Historically, it has not been easy to reach an agreement in the field on seemingly simple concepts such as what metrics are suitable to assess performance, the workload that should be used for stress tests, and even what performance means. Despite this, researchers were able to continuously improve the ability of computers to solve increasingly complex problems using an intuitive approach. This early success partially contributed to the lack of rigour in performance evaluation, and the delay of its formal introduction. Fortunately, some ground rules were set by some preliminary works on benchmarking 6448963 ; HowNotToLieWithStatistics ; Jacob1995NotesOC ; CharacterizingComputerPerformanceWithASingleDigit . These rules trained researchers in data aggregation using the most appropriate measures of central tendency, as well as they depicted common pitfalls when designing test suites. Moreover, they paved the way for the introduction of standardization through Standard Performance Evaluation Corporation (SPEC) and Transaction Processing Performance Council (TPC) HowToBuildABenchmark ; Eigenmann2001nl . Eventually, authors were able to compile a solid framework for the evaluation of performance for classical computers, offering a systematic approach to benchmarking Lilja2005o ; Jain1991-xw ; Kounev2020mh ; leboudec2010performance ; Feitelson2015 .
@@
The adoption of standardized benchmarks within a scientific community has been shown to have a positive impact on the overall scientific maturity of a discipline Sim_BenchmarkingToAdvanceResearch . Benchmarks can help to establish standards and best practices for research and development, as well as provide a basis for objective comparison between different systems. This can lead to more rigorous and reliable research practices, and ultimately advance the field as a whole. Nowadays, the quantum computing community faces similar challenges in measuring performance as those initially encountered by classical computing researchers. Even more, quantum computing is a nascent field, with numerous technology proposals and emerging issues further complicating the situation.
@@
Although quantum computers have not yet demonstrated a practical advantage over classical systems in real-world applications, this does not preclude the development of meaningful benchmarks. Moreover, there is theoretical evidence that quantum computers can offer significant advantages for certain problems Shor1994 . An analogous situation occurred during the early years of classical computing. During the early development of classical computing, benchmarks were primarily focused on evaluating low-level system, such as Whetstone WhetstoneBook and Dhrystone DhrystonePaper , rather than application-specific tasks. These benchmarks played a critical role in guiding hardware development and assessing system maturity.
@@
In the early stages of classical computing, the lack of standardized benchmarking rules allowed manufacturers to define their own benchmark, often leading to both unintentional and strategic biases. In some cases, manufacturers even aggressively optimized their compilers or CPUs to perform well on specific benchmarks Hennessy2012 . This trend contributed to the emergence of a ”benchmarking industry” driven more by marketing considerations than by objective performance evaluation.
@@
The field of quantum computing is similarly vulnerable. The lack of standardized benchmarking rules, together with the increasing competition among major industry players vying to outperform their rivals in the quantum computing race, might lead to a situation that resonates with Goodhart’s law goodhart1975problems : ”Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes”. In other words, the absence of a rigorously formulated and scrutinized approach to benchmarking in quantum computing may distort research priorities and hinder the development of truly scalable and reliable quantum processors. This bias occurs due to benchmarking models playing a crucial role on favouring what portion of the total ”behavior space” is captured SkadronChallengesInComputerArchitectureEvaluation .
@@
This is an already known challenge in the field of quantum computing osti_1775058 ; osti_1648785 ; osti_1639885 ; osti_1594677 ; quantumComputingSummitWhitepaperAtlanta , and there are initiatives such as the P7131 Project Authorization Request (PAR) proposal from the IEEE to standarize quantum computing performance hardware and software benchmarking StandrdForQuantumComputingPerformanceIEEE , or the proposals from IBM to define key characteristics of a quantum benchmark amico2023defining . Still, again, as the Goodhart’s law warned us, and the classic article from Dongarra et al. stated BlumeKohout2019 ; 6448963 : ”Although benchmarks are essential in performance evaluation, simple-minded application of them can produce misleading results. In fact, bad benchmarking can be worse than no benchmarking at all.” Therefore, we believe there is a clear need for a standardized benchmarking framework in the field of quantum computing to be able to assess how good the different attempts are at measuring performance.
@@
This article is organized as follows. In Sec. II, we provide an overview of the key elements of both classical processor benchmarks and metrics, emphasizing the fundamental properties they must exhibit to be considered good benchmarks and metrics. Additionally, we establish a common definition of various concepts used in the field of classical benchmarking. Finally, we review some examples of classical benchmarks and their limitations. Then, in Sec. III, we introduce some of the distinctive features of quantum computing devices that distinguish them from classical counterparts. We unify and define concepts to be used in the field of quantum benchmarking based on the definition of their classical counterpart. Furthermore, we review existing proposals for assessing the performance of a quantum computer, presenting the most relevant techniques, metrics, and benchmarks, and assessing which quality attributes these metrics and benchmarks fulfill. In Sec. IV, we propose general guidelines for any practitioner to follow when proposing a quantum benchmark. We also compile a series of open questions that need resolution on the path toward standardizing the comparison of quantum processor performance. Subsequently, we propose the creation of an organization dedicated to Standard Performance Evaluation for Quantum Computers (SPEQC). Finally, in Sec. V, we present our concluding remarks.
II What can we learn from Classical benchmarking?
@@
Originally, the word benchmark used to refer to ”a mark on a workbench used to compare the lengths of pieces so as to determine whether one was longer or shorter than desired” specglossary . Even though the term has clearly evolved, it retains the intuitive idea of comparing one object against another one to measure its properties. Hence, it is not a surprise that the concept of computer benchmark PP is based on this very same idea, and it is usually defined as ”a test, or set of tests, designed to compare the performance of one computer system against the performance of others” specglossary .
@@
A natural question that then arises is what we mean by performance. Performance is typically conceived in the classical computing scope as the useful work that a computer can get done in a given amount of time, and/or with a set of given resources. Indeed, in the past, it was common to hold the premise that ”the time required to perform a specified amount of computation is the ultimate measure of computer performance” CharacterizingComputerPerformanceWithASingleDigit . For classical computers, there have been historically two major approaches to the performance analysis problem: (1) Experimenting and measuring the actual system, and (2) Experimenting with a model of the system. For the latter, for any kind of performance evaluation not limited to computing devices you can use a physical model or a mathematical model. For computational performance assessment, we prefer to build models that can represent the system in terms of quantitative and logical relations. These models can later be either solved using mathematical tooling to obtain an analytical solution, or simulated Zeigler2018fg ; SimulationModelingAndAnalysisLawAverill . Even though in this work we will focus on the measurements of experiments with the actual systems, it is possible that we will see other approaches involving system modeling in the future. In fact, model-based approaches have been used by the quantum community for the assessment of quantum networks Wu2021 ; Coopmans2021 and the characterization of both incoherent and coherent errors in quantum gates Tripathi . Furthermore, other event-based models of quantum phenomena have also been published DeRaedt2005 ; Willsch2020 .
@@
In order to accurately evaluate the performance of devices, it is important to consider their various characteristics. Based on past attempts, it has become evident that a comprehensive assessment requires a holistic approach that takes into account multiple factors CharacterizingComputerPerformanceWithASingleDigit . Furthermore, the tests used in such assessments must meet certain standards of fairness to ensure that the results are not biased. Therefore, a rigorous and well-defined methodology, that is specifically tailored to the device being evaluated, is required. Failure to do so could result in inaccurate or misleading conclusions.
@@
II.1 Characteristics of a good benchmark
Before delving into the characteristics that define a good device/processor benchmark, and the importance of considering these factors when creating and using them, it is important to establish a common definition of the term. Although there are various interpretations of the term, we will agree with the definition of the term proposed by the Standard Performance Evaluation Corporation (SPEC) specglossary :
@@ II.1 Characteristics of a good benchmark
This terminology has been widely accepted in the field and provides a comprehensive definition of what a benchmark should involve. It is worth noting that other sources have also proposed alternative definitions for the term Lilja2005o ; SkadronChallengesInComputerArchitectureEvaluation ; Kounev2020mh ; HowToBuildABenchmark ; leboudec2010performance ; Jain1991-xw . We would like to remark that a benchmark is chosen in order to test the performance with respect to a given test suite, i.e. an application or set of applications. A better performance for this task does not necessarily mean a better performance for other tasks.
According to the definition, the crux of creating a high-quality benchmark rests in selecting appropriate tests that precisely reflect the types of workloads the system is expected to encounter, or that measure the anticipated behavior of the system during specific tasks relevant to the assessor. In other words, the tests should be representative of the performance of the system in real-world scenarios (i.e. the portion of the behavior space that is considered relevant) and should provide a realistic measure of the capabilities of the system related to the task. On classical benchmarking, researchers have been deeply studying how to model the workload so that it is representative of the behavior it models. Note that this definition implies that the workload is context-agnostic, but that is not the case. A workload is only a good fit for a problem when restricted to a certain context (e.g., imagine a workload given by a sequence of operations. If the operations can be executed in parallel by the device, or if the performance metric is not correlated to the devices capacity of solving operations, the workload might not be useful) Feitelson2015 . For classical devices, special detail has been paid to understanding the main contributing factors that need to be considered when modelling a workload, and multiple examples can be found in the literature ComputerResearchWorkload ; ArtificialWorkloadDesign . Indeed, in Ref. Jain1991-xw , R. Jain explicitly states that the discussion on selecting the most appropriate workload is the deepest pitfall into which a performance analysis project may fall. Alternatively, in a simplistic manner, we can state that the workload is ultimately the input for the tests, and hence the old computer science classic garbage in, garbage out (GIGO) principle applies.
@@
To be able to build a good benchmark, we introduce below a list of quality traits to ensure that it is comprehensive, accurate, and representative of system performance. These attributes are not intended as strict definitions, but rather allow for a degree of flexibility, making them applicable across a wide range of contexts. This approach aligns with the benchmarking principles proposed by Kounev Kounev2020mh and also reflect the most desirable characteristics of benchmarks proposed by other authors in the field HowToBuildABenchmark ; GustafsonAndSnell_HINT ; Henning_SpecCpu2000 ; Huppler_TheArtOfBuildingGoodBenchmark ; Sim_BenchmarkingToAdvanceResearch .
@@
In particular, let us start by considering relevance, the first key attribute of an ideal benchmark. Crafting a benchmark that correlates with the performance of the system on a specific domain poses a considerable challenge, as it often results in limited applicability beyond that particular area. Conversely, benchmarks designed with the intention of broader applicability tend to sacrifice the depth of meaningfulness in a specific scenario. Thus, achieving the optimal balance of relevance in benchmark design requires careful deliberation and trade-offs, ensuring that the benchmarks scope aligns with the intended application domain while maintaining a meaningful level of specificity Huppler_TheArtOfBuildingGoodBenchmark .
@@
As we have introduced, fairness can be substantially influenced by the conditions before executing the benchmark. Even more, fairness requires ensuring that multiple configuration rules can be applied when running tests while ensuring a common ground is established to allow comparisons. To facilitate fair comparisons, on classical computing a mechanism known as peak and base metrics were introduced by SPEC CPU2017 cpu2017 . Base metrics require that the test programs, constituting the building modules of a test suite in a given language, are compiled using the same rules (i.e. compilation flags), in the same order. Also, these metrics are mandatory for the generation of the benchmark report. On the other hand, optional peak metrics can be included allowing for greater flexibility, enabling the usage of different compilation rules for each benchmark, as well as feedback-directed optimization.
@@
Even a meticulously designed benchmark incorporating the aforementioned key features is susceptible to manipulation by different agents with conflicting goals. As an illustrative example, in Ref. Jain1991-xw , R. Jain presents a thought-provoking discussion on strategies employed to gain an advantage in what he terms the ratio game. A ratio game is a technique that involves using ratios with incomparable bases and skillfully combining them to benefit oneself. This is a common scenario for benchmarking, whenever we want to assess performance among distinct devices, normalizing against the performance of a reference device. For this, even when the geometric mean is properly used to average the normalized values HowNotToLieWithStatistics , it is possible to derive rules for beating the ratio game. Upon analysis, three rules emerge regarding the ratio game:
@@
    Even when one system exhibits better performance than another on all benchmarks, the selection of an appropriate base P can be manipulated to showcase a more favorable relative performance.
@@
These statements highlight the susceptibility of benchmarks to manipulation, emphasizing the need for cautious interpretation and critical evaluation when drawing conclusions based on the results from ratio games. Other researchers such as Mashey have also call out the difficulties of aggregating data, and the controversy of using one Pythagorean mean over the others, as well as stated the common agreement that ”performance is not a single number” Mashey2004 . Nonetheless, even though there is agreement on the fact that a single number is delusional, there are still discussions to determine what is the preferred metric, sparking debates over which numerical representations are deemed more optimal in each context.
@@
Finally, it is important to remember that while a good benchmark can provide a reliable and accurate measure of system performance, it cannot solve the problem of evaluating conflicting aspects of a computer device. Instead, it can help to properly balance the trade-offs for a particular objective and influence its strengths and weaknesses. It is not possible for a single benchmark to excel in all areas, and there is no silver bullet to solve this problem SkadronChallengesInComputerArchitectureEvaluation . Therefore, multiple workloads and benchmarks are necessary to comprehensively evaluate the performance of a system. Each benchmark may focus on different aspects of performance and may use different metrics and methodologies. At best, we can obtain a set of benchmarks that will be representative of the behavior space we want to model and measure. By combining the results of these multiple benchmarks, it would be possible to obtain a comprehensive and accurate picture of the system. Ultimately, the choice of benchmark(s) should depend on the specific use case and the objectives of the evaluation, as the relevance requirement already stated.
II.2 What types of benchmarks are there?

### v4

@@ Abstract
keywords:
quantum computing, quantum computer performance, quantum benchmarking
1 Introduction
@@
Proper performance analysis of devices is key to providing information to potentially interested users, allowing fair comparisons among different devices. Additionally, manufacturers use these insights to keep improving their systems. The measurement of performance for classical computer systems has traditionally been controversial. Historically, it has not been easy to reach an agreement in the field on seemingly simple concepts such as what metrics are suitable to assess performance, the workload that should be used for stress tests, and even what performance means. Despite this, researchers were able to continuously improve the ability of computers to solve increasingly complex problems using an intuitive approach. This early success partially contributed to the lack of rigour in performance evaluation, and the delay of its formal introduction. Fortunately, some ground rules were set by some preliminary works on benchmarking [1, 2, 3, 4]. These rules trained researchers in data aggregation using the most appropriate measures of central tendency, as well as they depicted common pitfalls when designing test suites. Moreover, they paved the way for the introduction of standardization through Standard Performance Evaluation Corporation (SPEC) and Transaction Processing Performance Council (TPC) [5, 6]. Eventually, authors were able to compile a solid framework for the evaluation of performance for classical computers, offering a systematic approach to benchmarking [7, 8, 9, 10, 11].
@@
The adoption of standardized benchmarks within a scientific community has been shown to have a positive impact on the overall scientific maturity of a discipline [12]. Benchmarks can help to establish standards and best practices for research and development, as well as provide a basis for objective comparison between different systems. This can lead to more rigorous and reliable research practices, and ultimately advance the field as a whole. Nowadays, the quantum computing community faces similar challenges in measuring performance as those initially encountered by classical computing researchers. Even more, quantum computing is a nascent field, with numerous technology proposals and emerging issues further complicating the situation.
@@
Although quantum computers have not yet demonstrated a practical advantage over classical systems in real-world applications, this does not preclude the development of meaningful benchmarks. Moreover, there is theoretical evidence that quantum computers can offer significant advantages for certain problems [13]. An analogous situation occurred during the early years of classical computing. During the early development of classical computing, benchmarks were primarily focused on evaluating low-level system, such as Whetstone [14] and Dhrystone [15], rather than application-specific tasks. These benchmarks played a critical role in guiding hardware development and assessing system maturity.
@@
In the early stages of classical computing, the lack of standardized benchmarking rules allowed manufacturers to define their own benchmark, often leading to both unintentional and strategic biases. In some cases, manufacturers even aggressively optimized their compilers or CPUs to perform well on specific benchmarks [16]. This trend contributed to the emergence of a “benchmarking industry” driven more by marketing considerations than by objective performance evaluation.
@@
The field of quantum computing is similarly vulnerable. The lack of standardized benchmarking rules, together with the increasing competition among major industry players vying to outperform their rivals in the quantum computing race, might lead to a situation that resonates with Goodhart’s law [17]: “Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes”. In other words, the absence of a rigorously formulated and scrutinized approach to benchmarking in quantum computing may distort research priorities and hinder the development of truly scalable and reliable quantum processors. This bias occurs due to benchmarking models playing a crucial role on favouring what portion of the total “behavior space” is captured [18].
@@
This is an already known challenge in the field of quantum computing [19, 20, 21, 22, 23], and there are initiatives such as the P7131 Project Authorization Request (PAR) proposal from the IEEE to standarize quantum computing performance hardware and software benchmarking [24], or the proposals from IBM to define key characteristics of a quantum benchmark [25]. Still, again, as the Goodhart’s law warned us, and the classic article from Dongarra et al. stated [26, 1]: “Although benchmarks are essential in performance evaluation, simple-minded application of them can produce misleading results. In fact, bad benchmarking can be worse than no benchmarking at all.” Therefore, we believe there is a clear need for a standardized benchmarking framework in the field of quantum computing to be able to assess how good the different attempts are at measuring performance.
@@
This article is organized as follows. In Sec. 2, we provide an overview of the key elements of both classical processor benchmarks and metrics, emphasizing the fundamental properties they must exhibit to be considered good benchmarks and metrics. Additionally, we establish a common definition of various concepts used in the field of classical benchmarking. Finally, we review some examples of classical benchmarks and their limitations. Then, in Sec. 3, we introduce some of the distinctive features of quantum computing devices that distinguish them from classical counterparts. We unify and define concepts to be used in the field of quantum benchmarking based on the definition of their classical counterpart. Furthermore, we review existing proposals for assessing the performance of a quantum computer, presenting the most relevant techniques, metrics, and benchmarks, and assessing which quality attributes these metrics and benchmarks fulfill. In Sec. 4, we propose general guidelines for any practitioner to follow when proposing a quantum benchmark. We also compile a series of open questions that need resolution on the path toward standardizing the comparison of quantum processor performance. Subsequently, we propose the creation of an organization dedicated to Standard Performance Evaluation for Quantum Computers (SPEQC). Finally, in Sec. 5, we present our concluding remarks.
2 What can we learn from Classical benchmarking?
@@
Originally, the word benchmark used to refer to “a mark on a workbench used to compare the lengths of pieces so as to determine whether one was longer or shorter than desired” [27]. Even though the term has clearly evolved, it retains the intuitive idea of comparing one object against another one to measure its properties. Hence, it is not a surprise that the concept of computer benchmark [28] is based on this very same idea, and it is usually defined as “a test, or set of tests, designed to compare the performance of one computer system against the performance of others” [27].
@@
A natural question that then arises is what we mean by performance. Performance is typically conceived in the classical computing scope as the useful work that a computer can get done in a given amount of time, and/or with a set of given resources. Indeed, in the past, it was common to hold the premise that “the time required to perform a specified amount of computation is the ultimate measure of computer performance” [4]. For classical computers, there have been historically two major approaches to the performance analysis problem: (1) Experimenting and measuring the actual system, and (2) Experimenting with a model of the system. For the latter, for any kind of performance evaluation not limited to computing devices you can use a physical model or a mathematical model. For computational performance assessment, we prefer to build models that can represent the system in terms of quantitative and logical relations. These models can later be either solved using mathematical tooling to obtain an analytical solution, or simulated [29, 30]. Even though in this work we will focus on the measurements of experiments with the actual systems, it is possible that we will see other approaches involving system modeling in the future. In fact, model-based approaches have been used by the quantum community for the assessment of quantum networks [31, 32] and the characterization of both incoherent and coherent errors in quantum gates [33]. Furthermore, other event-based models of quantum phenomena have also been published [34, 35].
@@
In order to accurately evaluate the performance of devices, it is important to consider their various characteristics. Based on past attempts, it has become evident that a comprehensive assessment requires a holistic approach that takes into account multiple factors [4]. Furthermore, the tests used in such assessments must meet certain standards of fairness to ensure that the results are not biased. Therefore, a rigorous and well-defined methodology, that is specifically tailored to the device being evaluated, is required. Failure to do so could result in inaccurate or misleading conclusions.
@@
2.1 Characteristics of a good benchmark
Before delving into the characteristics that define a good device/processor benchmark, and the importance of considering these factors when creating and using them, it is important to establish a common definition of the term. Although there are various interpretations of the term, we will agree with the definition of the term proposed by the Standard Performance Evaluation Corporation (SPEC) [27]:
@@ 2.1 Characteristics of a good benchmark
This terminology has been widely accepted in the field and provides a comprehensive definition of what a benchmark should involve. It is worth noting that other sources have also proposed alternative definitions for the term [7, 18, 9, 5, 10, 8]. We would like to remark that a benchmark is chosen in order to test the performance with respect to a given test suite, i.e. an application or set of applications. A better performance for this task does not necessarily mean a better performance for other tasks.
According to the definition, the crux of creating a high-quality benchmark rests in selecting appropriate tests that precisely reflect the types of workloads the system is expected to encounter, or that measure the anticipated behavior of the system during specific tasks relevant to the assessor. In other words, the tests should be representative of the performance of the system in real-world scenarios (i.e. the portion of the behavior space that is considered relevant) and should provide a realistic measure of the capabilities of the system related to the task. On classical benchmarking, researchers have been deeply studying how to model the workload so that it is representative of the behavior it models. Note that this definition implies that the workload is context-agnostic, but that is not the case. A workload is only a good fit for a problem when restricted to a certain context (e.g., imagine a workload given by a sequence of operations. If the operations can be executed in parallel by the device, or if the performance metric is not correlated to the devices capacity of solving operations, the workload might not be useful) [11]. For classical devices, special detail has been paid to understanding the main contributing factors that need to be considered when modelling a workload, and multiple examples can be found in the literature [36, 37]. Indeed, in Ref. [8], R. Jain explicitly states that the discussion on selecting the most appropriate workload is the deepest pitfall into which a performance analysis project may fall. Alternatively, in a simplistic manner, we can state that the workload is ultimately the input for the tests, and hence the old computer science classic garbage in, garbage out (GIGO) principle applies.
@@
To be able to build a good benchmark, we introduce below a list of quality traits to ensure that it is comprehensive, accurate, and representative of system performance. These attributes are not intended as strict definitions, but rather allow for a degree of flexibility, making them applicable across a wide range of contexts. This approach aligns with the benchmarking principles proposed by Kounev [9] and also reflect the most desirable characteristics of benchmarks proposed by other authors in the field [5, 38, 39, 40, 12].
@@
In particular, let us start by considering relevance, the first key attribute of an ideal benchmark. Crafting a benchmark that correlates with the performance of the system on a specific domain poses a considerable challenge, as it often results in limited applicability beyond that particular area. Conversely, benchmarks designed with the intention of broader applicability tend to sacrifice the depth of meaningfulness in a specific scenario. Thus, achieving the optimal balance of relevance in benchmark design requires careful deliberation and trade-offs, ensuring that the benchmarks scope aligns with the intended application domain while maintaining a meaningful level of specificity [40].
@@
As we have introduced, fairness can be substantially influenced by the conditions before executing the benchmark. Even more, fairness requires ensuring that multiple configuration rules can be applied when running tests while ensuring a common ground is established to allow comparisons. To facilitate fair comparisons, on classical computing a mechanism known as peak and base metrics were introduced by SPEC CPU2017 [41]. Base metrics require that the test programs, constituting the building modules of a test suite in a given language, are compiled using the same rules (i.e. compilation flags), in the same order. Also, these metrics are mandatory for the generation of the benchmark report. On the other hand, optional peak metrics can be included allowing for greater flexibility, enabling the usage of different compilation rules for each benchmark, as well as feedback-directed optimization.
@@
Even a meticulously designed benchmark incorporating the aforementioned key features is susceptible to manipulation by different agents with conflicting goals. As an illustrative example, in Ref. [8], R. Jain presents a thought-provoking discussion on strategies employed to gain an advantage in what he terms the ratio game. A ratio game is a technique that involves using ratios with incomparable bases and skillfully combining them to benefit oneself. This is a common scenario for benchmarking, whenever we want to assess performance among distinct devices, normalizing against the performance of a reference device. For this, even when the geometric mean is properly used to average the normalized values [2], it is possible to derive rules for beating the ratio game. Upon analysis, three rules emerge regarding the ratio game:
@@
    Even when one system exhibits better performance than another on all benchmarks, the selection of an appropriate base [42] can be manipulated to showcase a more favorable relative performance.
@@
These statements highlight the susceptibility of benchmarks to manipulation, emphasizing the need for cautious interpretation and critical evaluation when drawing conclusions based on the results from ratio games. Other researchers such as Mashey have also call out the difficulties of aggregating data, and the controversy of using one Pythagorean mean over the others, as well as stated the common agreement that “performance is not a single number” [43]. Nonetheless, even though there is agreement on the fact that a single number is delusional, there are still discussions to determine what is the preferred metric, sparking debates over which numerical representations are deemed more optimal in each context.
@@
Finally, it is important to remember that while a good benchmark can provide a reliable and accurate measure of system performance, it cannot solve the problem of evaluating conflicting aspects of a computer device. Instead, it can help to properly balance the trade-offs for a particular objective and influence its strengths and weaknesses. It is not possible for a single benchmark to excel in all areas, and there is no silver bullet to solve this problem [18]. Therefore, multiple workloads and benchmarks are necessary to comprehensively evaluate the performance of a system. Each benchmark may focus on different aspects of performance and may use different metrics and methodologies. At best, we can obtain a set of benchmarks that will be representative of the behavior space we want to model and measure. By combining the results of these multiple benchmarks, it would be possible to obtain a comprehensive and accurate picture of the system. Ultimately, the choice of benchmark(s) should depend on the specific use case and the objectives of the evaluation, as the relevance requirement already stated.
2.2 What types of benchmarks are there?